3. The pH of a 0.345 M weak base solution is 9.39. 50.0 mL of the weak base are titrated with 0.425 M HCI. a. Calculate the Kb of the weak base. b. Calculate the Ka of the weak base. c. d. Consult a Chemistry text and determine the identity of the weak base. Calculate the volume of HCl required to reach the equivalence point.
Answer #1

QUESTION 3

a) First we calculate the OH- concentration using the following equation.

14-pH=-log[OH^-]

14-9.39=-log[OH^-]

[OH^-]=2.45*10^{-5}M

Consider the following equilibrium between the base B and water.

B+H_2O\rightleftharpoons BH^++OH^-

Note that the concentrations of BH+ and OH- are equal. The equation for Kb can be written as,

K_b=\frac{[BH^+][OH^-]}{[B]}=\frac{[OH^-]^{2}}{[B]}

Substituting the concentrations,

K_b=\frac{(2.45*10^{-5}M)^{2}}{0.345M-2.45*10^{-5}M}

We assume that 0.345M-2.45*10-5M is approximately equal to 0.345M.

K_b=\frac{(2.45*10^{-5}M)^{2}}{0.345M}=1.74*10^{-9}M

b) Consider the above equilibrium,

B+H_2O\rightleftharpoons BH^++OH^-

BH+ can dissociate to give,

BH^+\rightleftharpoons B+H^+

Therefore, the Ka equation can be given as,

K_a=\frac{[B][H^+]}{[BH^+]}

For a conjugate acid-base pair, we can write,

K_a*K_b=K_w

K_a=\frac{K_w}{K_b}

K_a=\frac{10^{-14}}{1.74*10^{-9}}=5.75*10^{-6}

c) Online sources state that the Kb value of pyridine is 1.7*10-9. Thus the identity is most likely to be pyridine.

d) The following equation can be used to determine the volume required to reach the equivalence point.

C_{1}V_{1}=C_{2}V_{2}

C_{HCl}V_{HCl}=C_{base}V_{base}

0.425M*V_{HCl}=0.345M*50.0ml

V_{HCl}=40.588ml=40.6ml